# Stochastic gravitational-wave background search using data from five pulsar timing arrays

### 23 Mar 2026 v2

@@ Abstract
Using public pulse time-of-arrival data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB). This analysis is more sensitive than previous individual PTA searches because the combined 121-pulsar dataset is about four times larger than any single PTA’s. For pulsars observed by multiple PTAs, we employ a new “direct combination” method to merge their astrophysical models and data. This avoids the challenge of reconciling different PTA timing models to obtain a single “best” model. A central result of our analysis is the posterior distribution of the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency. While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and noise-marginalized false-alarm probabilities ($p$-values) for three detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection. The inter-pulsar timing-residual correlation, reconstructed as a function of angle $\theta$ between the pulsar lines of sight, matches the Hellings and Downs (HD) prediction.
@@
Aside from this, our analysis is standard. We use the same tools and methods as the individual PTAs, as implemented in Enterprise [24] and in the next-generation PTA data-analysis package Discovery [25]. Both are based on parameterized likelihood models, and employ Markov chain Monte Carlo (MCMC) methods to produce posterior samples (here called “posteriors”) which approximate the inferred posterior probability distributions of the model parameters. Detection confidence is assessed using both Bayesian and frequentist methods.
Note: additional tables and plots may be found in the Supplementary Material (SM).
@@
– The null hypothesis is called “common uncorrelated red noise” and is denoted by CURN. Again, the timing residuals are assumed to be drawn from a zero-mean Gaussian process, but this time entirely characterized by the covariance matrix ${\bf C_{0}}={\rm diag}\,\bf C$. Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis. The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the inter-pulsar HD correlations [28], hence the name “uncorrelated”. It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms. Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR. Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2]. CURN is a more “conservative” null hypothesis than that of “no GWs”.
Using Discovery [25] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models. These are sets of inferred parameter values most consistent with the data and priors, thereby characterizing the posterior distributions of the model parameters.
@@
One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$. We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to take the value $13/3$. The latter would be expected from a population of SMBHBs in circular orbits, provided that GWs dominate the energy loss at nHz frequencies.
One way to assess detection confidence is via frequentist detection statistics [28]. We present results for three different statistics that are used within the PTA community. All three statistics $D$ are quadratic functions of the timing residuals, with weights determined by $\bf C$ and $\bf C_{0}$.
Each posterior corresponds to a different model universe, with its own amplitudes/exponents for the different noise sources, and thus its own covariance matrices $\bf C$ and $\bf C_{0}$. Hence, the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next. To assess detection confidence, we present the posterior distributions of these $p$-values. In the literature, its mean is called the “posterior predictive $p$-value” [29], similar in spirit to the “noise-marginalized optimal statistic” [30].
Some past work uses ad hoc methods for assessing the false-alarm probability $p$. “Sky scrambling” and “phase shifting” were intended to “remove any signal from the data”. The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [31, 32]. The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [32]. However, neither of these methods properly reproduces the null distribution of $D$ [33]. For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
We determine the false-alarm probability $p$ using a statistically rigorous method [34, 35, 33, 28]. Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [34, 33, 28]; the probability that it exceeds the observed value in each of our posterior samples is found analytically.
To aid interpretation, we also report the $p$-values as “equivalent $q\sigma$” levels: $q>0$ is chosen so that a random draw from a zero-mean Gaussian with variance $\sigma^{2}$ has probability $p$ of exceeding $q\sigma$. This “one-sided $p$ value” satisfies $2p=\mathrm{erfc}(q/\sqrt{2})$. For example, $1\sigma$ corresponds to $p=16\%$, $2\sigma$ corresponds to $p=2.3\%$, $3\sigma$ corresponds to $p=0.14\%$, $4\sigma$ corresponds to $p=3.2\times 10^{-5}$, and $5\sigma$ corresponds to $p=2.9\times 10^{-7}$. The latter, a false-alarm probability of less than one per three million, is the traditional threshold for confident detection [12].
@@
The data we analyze consist of $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ times of arrival (TOAs): the complete set of observations from 25 EPTA pulsars, 14 InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 32 PPTA pulsars. We merge these using the direct combination method [21, 22]. If not already present, rotation counts are added to the .tim files, which contain pulse TOAs and uncertainties. The .par files, containing astrophysical models and detector-specific parameters, are modified one pulsar at a time (the first step of “manual” data combination), as follows.
  1. 1.
  2. 2.
  3. 3.
    If the pulsar’s .par files use both TCB and TDB time conventions, then quantities whose units contain time are converted to TDB with Tempo2 [41] tools. See SM Sec. III.
  4. 4.
  5. 5.
The direct combination method [21, 22] adopts one set of parameter values $\beta_{0}$ for each pulsar, item 5 above. It works because each pulsar has exact but unknown values $\beta$ for these (for example, for the sky position). They enter the analysis via the nominal value $\beta_{0}$ [1, Eq. 7.2] and via a “design matrix” $\mathbf{M}$} [1, Eq. 7.4], which is the rate of change of the pulse arrival-time with respect to variations in $\beta$, evaluated at $\beta_{0}$. In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} at the same nominal $\beta_{0}$ for a given pulsar introduces negligible errors. The subsequent marginalization over $\epsilon=\beta-\beta_{0}$ [1, Eq. 7.6] fully accounts for the uncertainties in $\beta$ and for the variations in its nominal value across PTAs. Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same across PTAs. We therefore adopt the physically motivated DMGP model [42], used by all PTAs except NG and InPTA, which use the DMX model. The design matrix $\mathbf{M}$} allows the effects of DM fluctuations on the timing residuals to be computed from the fixed-in-time DM timing residuals.
The next analysis step is single-pulsar noise modeling. These determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq. 7.27]. This characterizes the timing residual fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eqs. 7.2 and 7.26], modeled with 30 and 100 frequency bins respectively. Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
Single-pulsar noise modeling determines two or three parameters per pulsar and backend. The WN covariance matrix $\mathbf{N}$} is set by these parameters and by the .tim file contents. NG, PPTA, MPTA, and seven EPTA backends use the parameters $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and $\mathrm{ECORR}$} [39, Eq. 2]. InPTA and the remaining EPTA backends only use $\mathrm{EFAC}$} and $\mathrm{EQUAD}$} [43].
@@
The WN parameters for a given pulsar are obtained from all of its (.tim files) TOA data. These are analyzed using JAXopt [44], with priors taken from the Table 1 rows marked “S”; the amplitude/exponent priors are used for all power laws. For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model, while for $\gamma>1$ the $f\to\infty$ divergences are absent [45, 46]. Note that $\mathrm{EQUAD}$} has the units used in Tempo [47], Tempo2 [41], and PINT [48], and is written as $\rm T2EQUAD$. The output is a “WN dictionary” of best-fit $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and (potentially) $\mathrm{ECORR}$} values, used later to obtain $\mathbf{N}$} [39, Eq. 2].
For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA analysis values. This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more reliable estimates of the WN parameters. In SM Sec. I, we show that our conclusions are robust: they are not changed much by dropping TOAs from backends with the largest $\mathrm{EFAC}$} corrections.
The final stage is multi-pulsar analysis, following the procedures described in [1, 9]. We use both Enterprise [24] and the next generation PTA data-analysis package, Discovery, which is optimized via JAX for graphics processing units (GPU) [49]. The use of Discovery has significantly sped up our analysis.
Within the Discovery framework, we use the WN dictionary plus the SGWB priors listed in Tab. 1 with an “F” in the final column. As described above, we consider two priors for the SGWB power-law power-spectrum exponent $\gamma_{gw}$. One is the uniform range given in Tab. 1, the other is fixed to $\gamma_{gw}=13/3$. Our analysis uses the Hamiltonian No-U-Turn Sampler (NUTS) [50, 51], as implemented in NumPyro [52, 53, 54]. NUTS is the default for Discovery [25], which provides the log-likelihood gradients needed by Hamiltonian samplers. It generates samples that are much more independent than the parallel tempering MCMC sampler [55] previously used by the individual PTAs, and is also much faster.
@@
Results of the five PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1). These are shown in Fig. 2. The amplitude posterior is sharply peaked away from zero, which is consistent with (or evidence for) the presence of an SGWB. As expected, the posteriors are very similar for the two hypotheses. This is because, if there are $\lesssim 1500$ pulsars (see [56, Fig. 10]), most of the SGWB information is provided by auto- rather than cross-correlations, and by construction, CURN includes these autocorrelations.
Model $\log_{10}A_{gw}$ $\gamma_{gw}$
CURN with free $\gamma_{gw}\qquad$ $14.37^{+0.11}_{-0.12}\qquad$ $3.63^{+0.27}_{-0.26}$
CURN with $\gamma_{gw}=13/3\qquad$ $14.69^{+0.04}_{-0.04}\qquad$
HD with free $\gamma_{gw}\qquad$ $14.45^{+0.13}_{-0.15}\qquad$ $3.76^{+0.31}_{-0.29}$
HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.04}_{-0.04}\qquad$
@@
Table 2: Median posterior values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
@@
These posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6]. They are more constraining than the posteriors obtained by the individual PTAs: the uncertainty ellipses in the bottom-left plot of Fig. 2 are smaller (by about a factor of two) than those obtained by intersecting the individual PTA posteriors [57]. For example, the $\text{area}\equiv\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse for the HD hypothesis is 48% of that shown in [57, Fig. 1, right]. As explained earlier, this is as expected for an SGWB: joint analysis yields more evidence and is more constraining than naive intersection of the separate PTA results.
@@
The PTA community has established basic criteria for claiming a detection [12]. Among these, one must show that the data is consistent with the HD signal hypothesis, and very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis. The estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are very similar under both hypotheses.
@@
One way to assess significance is the frequentist approach: null-hypothesis testing [58] using a statistic. This computes the (noise-marginalized false-alarm) probability $p$ that signal-like data would occur under the null hypothesis. We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [28, Eq. (1)]). The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [30, 33, 28]. The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$ [28]. The NP minimum-variance (NPMV) statistic [28] adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
Why? Normally, the NP statistic would be preferred. However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise. In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them more robust. Details may be found in [28].
@@
The detection statistic values and their false-alarm probabilities $p$ depend on the observational data and (inferred) pulsar-noise models under the null (CURN) hypothesis. While the data are fixed, the models are uncertain and vary across posteriors, so we adopt a noise-marginalized approach [30, 29]. For each posterior-sample CURN universe, we compute the three statistics and evaluate their false-alarm probabilities using the corresponding generalized-$\chi^{2}$ distribution [34, 33]. We then form a histogram of the resulting $p$-values.
We first discuss the NPMV statistic, whose posterior $p$-value histogram is shown in Fig. 3. The mean/median false-alarm probabilities are $p=5.5\times 10^{-4}$$p=1.1\times 10^{-4}$ for free $\gamma_{gw}$ and $p=3.7\times 10^{-4}$$p=8.0\times 10^{-5}$ for fixed $\gamma_{gw}=13/3$. Respectively, these correspond to $3.3\sigma$$3.7\sigma$ and $3.4\sigma$$3.8\sigma$ equivalent Gaussian significances. The most likely range is shown in Fig. 3 as the band that contains 68% of the posterior points about the median. These correspond to significances in the range from $3.1\sigma$ to $4.3\sigma$ for free $\gamma_{gw}$ and $3.3\sigma$ to $4.3\sigma$ for $\gamma_{gw}=13/3$.
@@
The corresponding histograms for the NP and OS false-alarm probabilities $p$ are similar (see SM Sec. II). We list key results for all three statistics in Table 3. None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ detection threshold [12]. However, note that our analysis yields a traditional OS detection significance of $\approx 4.3\sigma$, which is higher than the $\approx 3.2\sigma$ significance reported by NG [35]. This suggests that the “direct data combination” method is working as would be expected in the presence of an SGWB.
Another way to quantify detection confidence [12] is via the Bayes factor ${\mathrm{BF}}=E_{\rm{HD}}/E_{\rm{CURN}}$, the ratio of HD to CURN model evidences. Here, the evidence $E$ is the likelihood marginalized over the prior [59]; $E$ has units $({\rm time})^{-N}$ where $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ is the number of TOAs (see SM Sec. I Table 4). Evaluating these integrals is challenging because the parameter space has $4\times 121+2=486$ dimensions for $\gamma_{gw}$ free ($485$ for $\gamma_{gw}=13/3$). To compute the evidences we use generalized stepping-stone sampling [60, 61, 62], which broadens the sharply peaked likelihood using $K=8$ chains (K1=7K-1=7 steps) while restricting the integration region. This gives $\ln(E_{\rm{HD}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}504.39\pm 0.10$ and $\ln(E_{\rm{CURN}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}494.21\pm 0.09$, implying $\ln({\mathrm{BF}})=10.18\pm 0.13$ and ${\mathrm{BF}}=26{\kern 0.83298pt}000\pm 3{\kern 0.83298pt}000$. This is two orders-of-magnitude larger than the Bayes factor ${\mathrm{BF}}\approx 200$ reported by NG [9]. For a real Gaussian random variable, the corresponding significance [29, Eq. (8)] of $\sqrt{2\ln({\mathrm{BF}})}\,\sigma\approx 4.5\sigma$ is below the threshold for confident detection.
@@
An optimal method for “reconstructing” the HD correlation from the data has been derived [63] but is not yet fully implemented [46]. So, our reconstruction is based on the same method [56] previously used by EPTA [7] and NG [9], as implemented in [64]. Fig. 4 shows the result for the $20{\kern 0.83298pt}480$ free $\gamma_{gw}$ CURN posterior samples.
Our HD reconstruction uses the NG [9] choice of 15 angular bins. The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin has 484 pairs. For each posterior sample, the correlations of the pulsar pairs are weighted as in [64], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ per bin. For each bin, we then construct and average the $20{\kern 0.83298pt}480$ corresponding Gaussian probability distributions. The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked. Fig. 4 shows their means (close to medians) and standard deviations. A traditional $\chi^{2}$ goodness-of-fit test shows that the reconstruction matches the HD curve well: the reduced $\chi^{2}$ of $0.74$ is consistent with expected statistical fluctuations for 15 degrees of freedom
Conclusion — We search for an SGWB using public data from five PTAs. In effect, we repeat an established analysis, but with a larger dataset. This is enabled by the physically motivated step of adopting consistent astrophysical parameters and noise models for all pulsars.
Our results provide the strongest evidence to date that an SGWB contributes to observed pulsar pulse-arrival-time fluctuations. However, the statistical significance—–assessed using multiple frequentist and Bayesian methods—–remains too low to claim a detection.
Using existing analysis pipelines enables straightforward comparison with individual PTA results, with which our findings are consistent, but it has limitations. For example, these pipelines assume that for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent. Although this assumption is not strictly valid [63, see Conclusion], its impact may be negligible or correctable [45]. Likewise, our HD-curve reconstruction is suboptimal because implementation of an optimal method [63] is still in progress [46]. Another limitation is the lack of a hierarchical pulsar-noise model [65, 66], which could shift the inferred SGWB amplitude and spectral slope [43].

### 28 Apr 2026 v3

@@ Abstract
Using public pulse time-of-arrival data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB). This analysis is more sensitive than previous individual PTA searches because the combined 121-pulsar dataset is about four times larger than any single PTA’s. For pulsars observed by multiple PTAs, we employ a new “direct combination” method to merge their astrophysical models and data. This avoids the challenge of reconciling different PTA timing models to obtain a single “best” model. A central result of our analysis is the posterior distribution of the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency. While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance-assessed via a Bayesian odds ratio and noise-marginalized false-alarm probabilities ($p$-values) for three detection statistics-remains below the conventional $5\sigma$ threshold for a confident detection. The inter-pulsar timing-residual correlation, reconstructed as a function of angle $\theta$ between the pulsar lines of sight, matches the Hellings and Downs (HD) prediction.
@@
Aside from this, our analysis is standard. We use the same tools and methods as the individual PTAs, as implemented in Enterprise [24] and in the next-generation PTA data-analysis package Discovery [25]. Both use parameterized likelihood models and Markov chain Monte Carlo (MCMC) to generate posterior samples (“posteriors”) approximating the parameters’ posterior distributions.
Note: additional details are provided in End Matter (EM) and supporting tables and plots in the Supplementary Material (SM).
@@
– The null hypothesis is called “common uncorrelated red noise” and is denoted by CURN. Again, the timing residuals are assumed to be drawn from a zero-mean Gaussian process, but this time entirely characterized by the covariance matrix ${\bf C_{0}}={\rm diag}\,\bf C$. This zeros all elements where the row and column have different pulsars, see A. Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis. Using Discovery [25] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models. These are sets of inferred parameter values most consistent with the data and priors, thereby characterizing the posterior distributions of the model parameters.
@@
One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$. We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to $\gamma_{gw}=13/3$. The latter would be expected from a population of SMBHBs in circular orbits, provided that GWs dominate the energy loss at nHz frequencies.
One way to assess detection confidence is via frequentist detection statistics [28]. We consider three different statistics $D$ used within the PTA community. Each $D$ is a quadratic function of the timing residuals, with weights determined by $\bf C$ and $\bf C_{0}$. Because each posterior sample corresponds to a different noise model and SGWB—and thus different covariances—the resulting statistic values and associated false-alarm probabilities ($p$-values) vary across samples. We therefore characterize detection confidence using the posterior distribution of these $p$-values, whose mean corresponds to the posterior predictive $p$-value [29], analogous to the noise-marginalized optimal statistic [30].
Some past work uses ad hoc techniques to estimate the false-alarm probability $p$, see A. We compute $p$-values using a statistically rigorous analytic approach [31, 32, 33, 28]. Since each statistic $D$ is a quadratic form in the data, its null (CURN) distribution is a generalized $\chi^{2}$; the probability of exceeding the observed value is obtained analytically for each posterior sample.
To aid interpretation, we also report the $p$-values as “equivalent $q\sigma$” levels: $q>0$ is chosen so that a random draw from a zero-mean Gaussian with variance $\sigma^{2}$ has probability $p$ of exceeding $q\sigma$. This “one-sided $p$ value” satisfies $2p=\mathrm{erfc}(q/\sqrt{2})$. For example, $1\sigma$ corresponds to $p=16\%$, $3\sigma$ corresponds to $p=0.14\%$, and $5\sigma$ corresponds to $p=2.9\times 10^{-7}$. The latter, a false-alarm probability of less than one per three million, is the traditional threshold for confident detection [12].
@@
The data we analyze consist of $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ times of arrival (TOAs): the complete set of observations from 25 EPTA pulsars, 14 InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 32 PPTA pulsars. We merge these using the direct combination method [21, 22]. For each pulsar, this adopts one set of parameter values $\beta_{0}$, defining its nominal astrophysical model [1, Eq. 7.2]; for details see A and SM Sec. A.1. It works because each pulsar has exact but unknown values $\beta$ for these parameters (for example, for the sky position). They enter the analysis via the nominal value $\beta_{0}$ [1, Eq. 7.2] and via a “design matrix” $\mathbf{M}$} [1, Eq. 7.4], which is the rate of change of the pulse arrival-time with respect to variations in $\beta$, evaluated at $\beta_{0}$. In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} at the same nominal $\beta_{0}$ for a given pulsar introduces negligible errors. The subsequent marginalization over $\epsilon=\beta-\beta_{0}$ [1, Eq. 7.6] fully accounts for the uncertainties in $\beta$ and for the variations in its nominal value across PTAs. Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same across PTAs. We therefore adopt the physically motivated DMGP model [39], used by all PTAs except NG and InPTA, which use the DMX model. The design matrix $\mathbf{M}$} allows the effects of DM fluctuations on the timing residuals to be computed from the fixed-in-time DM timing residuals.
The next analysis step is single-pulsar noise modeling, which determines each pulsar’s WN covariance matrix $\mathbf{N}$} appearing in the Gaussian likelihood [1, Eq. 7.27] (see A). This modeling yields two or three parameters per pulsar and backend-$\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and (in some cases) $\mathrm{ECORR}$}. NG, PPTA, MPTA, and seven EPTA backends use all three parameters [37, Eq. 2], while InPTA and the remaining EPTA backends use only $\mathrm{EFAC}$} and $\mathrm{EQUAD}$} [40].
@@
The WN parameters for each pulsar are inferred from all of its .tim-file TOA data using JAXopt [41], with priors from Table 1 (rows marked “S”); the amplitude/exponent priors are applied to all power laws. For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model, while for $\gamma>1$ the $f\to\infty$ divergences are absent [42, 43]. The parameter $\mathrm{EQUAD}$} is expressed in the Tempo/Tempo2/PINT convention (T2EQUAD[44, 45, 46]. The output is a “WN dictionary” of best-fit parameters, from which $\mathbf{N}$} is then determined using the .tim file contents.
The final stage is multi-pulsar analysis, following the procedures described in [1, 9]. We use both Enterprise [24] and the next generation PTA data-analysis package, Discovery, which is optimized via JAX for graphics processing units [47]. The use of Discovery has significantly sped up our analysis.
Within the Discovery framework, we use the WN dictionary plus the SGWB priors listed in Tab. 1 with an “F” in the final column. As described above, we consider two priors for the SGWB power-law power-spectrum exponent $\gamma_{gw}$. One is the uniform range given in Tab. 1, the other is fixed to $\gamma_{gw}=13/3$. Our analysis uses the Discovery [25] default Hamiltonian No-U-Turn Sampler (NUTS) [48, 49], as implemented in NumPyro [50, 51, 52], see A.
@@
Results of the five PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1). These are shown in Fig. 2. The amplitude posterior is sharply peaked away from zero, which is consistent with (or evidence for) the presence of an SGWB. As expected, the posteriors are very similar for the two hypotheses. This is because, if there are $\lesssim 1500$ pulsars (see [53, Fig. 10]), most of the SGWB information is provided by auto- rather than cross-correlations, and by construction, CURN includes these autocorrelations.
Model $\quad\log_{10}A_{gw}$ $\gamma_{gw}$ $\quad\log_{10}\Omega_{gw}$
CURN ($\gamma_{gw}$ free) $\quad-14.37^{+0.11}_{-0.12}$ $\quad 3.63^{+0.27}_{-0.26}$ $\quad-9.00^{+0.08}_{-0.09}$
CURN ($\gamma_{gw}$ fixed) $\quad-14.69^{+0.04}_{-0.04}$ $\quad 13/3$ $\quad-8.92^{+0.08}_{-0.08}$
HD ($\gamma_{gw}$ free) $\quad-14.45^{+0.13}_{-0.15}$ $\quad 3.76^{+0.31}_{-0.29}$ $\quad-9.05^{+0.09}_{-0.09}$
HD ($\gamma_{gw}$ fixed) $\quad-14.72^{+0.04}_{-0.04}$ $\quad 13/3$ $\quad-8.98^{+0.08}_{-0.09}$
@@
Table 2: Median posterior values for $A_{gw}$ and $\gamma_{gw}$, their 68% containment intervals from Fig. 2, and $\Omega_{gw}(3\,\text{nHz})$.
@@
These posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6]. They are more constraining than the posteriors obtained by the individual PTAs: the uncertainty ellipses in the bottom-left plot of Fig. 2 are smaller (by about a factor of two) than those obtained by intersecting the individual PTA posteriors [54]. For example, the $\text{area}\equiv\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse for the HD hypothesis is 48% of that shown in [54, Fig. 1, right]. As explained earlier, this is as expected for an SGWB: joint analysis yields more evidence and is more constraining than naive intersection of the separate PTA results.
@@
The PTA community has established basic criteria for claiming a detection [12]. Among these, one must show that the data is consistent with the HD signal hypothesis, and very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis. The estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are very similar under both hypotheses (see Fig. 2).
@@
As previously described, one way to assess significance is via frequentist null-hypothesis testing [55] using a statistic. This computes the (noise-marginalized false-alarm) probability $p$ that signal-like data would occur under the null hypothesis. We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the column vector of timing residuals, and consider three different standard choices of the square filter matrix $\mathbf{Q}$} (see A and [28, Eq. (1)]).
@@
While the data are fixed, the inferred pulsar and SGWB noise models-and hence the CURN null-hypothesis distributions-vary across posterior samples. We therefore adopt a noise-marginalized approach [30, 29], computing each statistic and its false-alarm probability for every posterior realization using the corresponding generalized $\chi^{2}$ distribution [31, 33]. The resulting $p$-value distributions are shown in SM Sec. A.2.
@@
We summarize the results for all three statistics in Table 3. None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ detection threshold [12]. However, note that our analysis yields a traditional OS detection significance of $\approx 4.3\sigma$, which is higher than the $\approx 3.2\sigma$ significance reported by NG [32]. This suggests that the “direct data combination” method is working as would be expected in the presence of an SGWB.
Another way to quantify detection confidence [12] is via the Bayes factor ${\mathrm{BF}}=E_{\rm{HD}}/E_{\rm{CURN}}$, the ratio of HD to CURN model evidences. Evaluating these (see A) gives $\ln({\mathrm{BF}})=10.18\pm 0.13$ $\implies$ ${\mathrm{BF}}=26{\kern 0.83298pt}000\pm 3{\kern 0.83298pt}000$. This is two orders-of-magnitude larger than the ${\mathrm{BF}}\approx 200$ reported by NG [9]. For a real Gaussian random variable, the corresponding significance [29, Eq. (8)] of $\sqrt{2\ln({\mathrm{BF}})}\,\sigma\approx 4.5\sigma$ is below the threshold for confident detection.
@@
An optimal method for “reconstructing” the HD correlation from the data has been derived [56] but is not yet fully implemented [43]. So, our reconstruction is based on the same method [53] previously used by EPTA [7] and NG [9], as implemented in [57] (for further details, see A). Fig. 3 shows the result for the $20{\kern 0.83298pt}480$ free $\gamma_{gw}$ CURN posterior samples. A traditional $\chi^{2}$ goodness-of-fit test shows that the reconstruction matches the HD curve well: the reduced $\chi^{2}$ of $0.74$ is consistent with expected statistical fluctuations for 15 degrees of freedom
Conclusion - We search for an SGWB using public data from five PTAs. In effect, we repeat an established analysis, but with a larger dataset (see A). This is enabled by the physically motivated step of adopting a single astrophysical/noise model for each pulsar.
Our results provide the strongest evidence to date that an SGWB contributes to observed pulsar pulse-arrival-time fluctuations. However, the statistical significance—assessed using multiple frequentist and Bayesian methods—remains too low to claim a detection.
Using existing analysis pipelines enables straightforward comparison with individual PTA results, with which our findings are consistent, but it has limitations. For example, these pipelines assume that for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent. Although this assumption is not strictly valid [56, see Conclusion], its impact may be negligible or correctable [42]. Likewise, our HD-curve reconstruction is suboptimal because implementation of an optimal method [56] is still in progress [43]. Another limitation is the lack of a hierarchical pulsar-noise model [58, 59], which could shift the inferred SGWB amplitude and spectral slope [40].